The number of workers needed for a bank is shown below.

Monday Tuesday Wednesday Thursday Friday Saturday Sunday
4 5 6 3 8 9 2

Suppose those employees in the same bank are paid $150 per day the first four days and they work a day of overtime at a cost of $350. How should the bank schedule its employees?
Answer #1

Monday

Tuesday Wednesday Thursday Friday    Saturday Sunday
4 5 6 3 8 9 2
4*150 = 600 5*150 = 750 6*150=900 3*150=450 8*350 = 2800 9*350=3150 2*350 = 700

Since the cost of having employees on Friday(8*350 = 2800), Saturday(9*350=3150) and Sunday(2*350=700) is $6650 the bank should schedule the work in the week days when the price paid is $150 for each day and each dollar.

THE BANK SHOULD SPLIT THE WORKING HOUSE THE DURING MONDAY TO THURSDAY.

If each day we have ten employee each day at the bank it will called so.

Total cost if 10 each winter. 4000 Rupees and we will be left with

Monday

Tuesday Wednesday Thursday Friday    Saturday Sunday
10 10 10 10 Holiday Holiday Holiday
4 5 6